OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y

pooja195 (pooja195):

4x + 2y = 11 <---equation 1 x – 2 = – 2y <---equation 2 It wants you to use substitiuion so you need to isolate a vairble x – 2 = – 2y add 2 to both sides what are you left with?

OpenStudy (anonymous):

3-1/2 is my answer

OpenStudy (anonymous):

right?

pooja195 (pooja195):

You mean (3,-1/2) which is correct

OpenStudy (anonymous):

yea lol

pooja195 (pooja195):

yw

OpenStudy (anonymous):

thanks